OpenStudy (anonymous):

solve for x (x+2)^2=(5+2)^2

OpenStudy (anonymous):

is there x in second part..?

OpenStudy (anonymous):

whoops ya (5+2x)^2

OpenStudy (anonymous):

x+2=5+2x x=-3

OpenStudy (anonymous):

you are taking square root of both sides are you ok?

OpenStudy (anonymous):

there should be 2 answers to it

OpenStudy (anonymous):

-7/3 the second one

OpenStudy (anonymous):

i unfolded squares on a both sides & solve it as quadratic equation :)

OpenStudy (anonymous):

did u end up with 3x^2 +16x +21 and then use the quadratic formula

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

how would you solve it by completing the square

OpenStudy (anonymous):

i used quadratic formula: http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

:)